# Toeplitz Operators Defined between Köthe Spaces

### v1

@@ 4. Operators Defined by an Upper Triangular Toeplitz Matrix
Let $\Lambda_{\infty}(\alpha)$ be power series space of infinite type and $K(b_{n,k})$ be a nuclear Köthe space. $\widecheck{T}_{\theta}:\Lambda_{\infty}(\alpha)\to K(b_{n,k})$ is compact if and only if $\theta\in(\Lambda_{\infty}(\alpha))^{\prime}$ and the following condition holds:
$\quad\quad\exists m\in\mathbb{N}\quad\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad\quad\quad b_{n,k}\leq Ce^{m\alpha_{n}}\quad\quad\forall n\in\mathbb{N}$ (4.2)
@@
This gives us that there exists a $C_{1}>0$ such that
$\|\widecheck{T}_{\theta}e_{n}\|=\sup_{1\leq j\leq n}|\theta_{n-j}|b_{j,k}\leq C_{1}\|e_{n}\|_{m}=C_{1}e^{m\alpha_{n}}$
By assuming $\theta_{0}\neq 0$, we can write that for every $k\in\mathbb{N}$ there exists a $C>0$ such that
$b_{n,k}\leq Ce^{k\alpha_{n}}$
for all $n\in\mathbb{N}$. This says that 4.2 is satisfied. For the converse, assume $\theta\in(\Lambda_{\infty}(\alpha))^{\prime}$ and 4.2 holds. By the same calculation of the proof of the Theorem 4.2, we can write that there exists a $m\in\mathbb{N}$ such that for all $k\in\mathbb{N}$
@@
Let $\Lambda_{\infty}(\alpha)$ be power series space of infinite type and $K(b_{n,k})$ be a Köthe space. Assume that there exists a non-decreasing function $S:\mathbb{N}\to\mathbb{N}$ such that the following condition holds
$\quad\quad\quad\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad\quad\quad b_{n,k}\leq Ce^{S(k)\alpha_{n}}\quad\quad\quad\forall n\in\mathbb{N}.$ (4.3)
@@
Let assume $\theta\in(\Lambda_{\infty}(\alpha))^{\prime}$. Then there exists some $m_{1}\in\mathbb{N}$ and $C_{1}>0$ such that
$|\theta_{n-1}|\leq C_{1}e^{m_{1}\alpha_{n}}\quad\quad\forall n\in\mathbb{N}.$
For every $k\in\mathbb{N}$ satisfying $S(k)\geq m_{1}$, there exists $C_{2}>0$ such that
$\begin{split}\|\widecheck{T}_{\theta}e_{n}\|_{k}&=\sup_{1\leq j\leq n}|\theta_{n-j}|b_{j,k}\\ &\leq C_{1}\sup_{1\leq j\leq n}e^{m_{1}\alpha_{n-j+1}}b_{j,k}\\ &\leq C_{2}\sup_{1\leq j\leq n}e^{m_{1}\alpha_{n}}e^{S(k)\alpha_{j}}\\ &\leq C_{2}e^{(m_{1}+S(k))\alpha_{n}}\leq C_{2}e^{2S(k)\alpha_{n}}=\|e_{n}\|_{2S(k)}\end{split}$
@@
$\|\widecheck{T}_{\theta}e_{n}\|_{k}\leq C_{2}\|e_{n}\|_{2S(k)}\hskip 43.36243pt$
for all $k$ satisfying $S(k)\geq m_{1}$, that is, the family $\widecheck{\mathcal{A}}_{\infty}$ is 2S-tame. ∎
For the converse of Proposition 4.3, we again need the condition 3.4 on the sequence $\alpha$ when $K(a_{n,k})$ is power series space of finite type $\Lambda_{1}(\alpha)$.
@@
Let $\alpha=\left(\alpha_{n}\right)_{n\in\mathbb{N}}$ be a non-negative increasing sequence with $\lim_{n\rightarrow\infty}\alpha_{n}=\infty$. Assume that $\alpha$ satisfies the condition 3.4. $\widecheck{T}_{\theta}:\Lambda_{1}(\alpha)\to K(b_{n,k})$ is continuous if and only if $\theta\in(\Lambda_{1}(\alpha))^{\prime}$ and the following condition holds
$\quad\quad\forall k\in\mathbb{N}\quad\exists m\in\mathbb{N},C>0\quad\quad\quad\quad\quad b_{n,k}\leq Ce^{-\frac{1}{m}\alpha_{n}}\quad\quad\forall n\geq k.$ (4.4)
@@
The sufficiency part comes from Proposition 4.1. For the necessary part, assume $\theta\in(\Lambda_{1}(\alpha))^{\prime}$. Then there exists some $m_{1}\in\mathbb{N}$ and $C_{1}>0$ such that
$|\theta_{n-1}|\leq C_{1}e^{-\frac{1}{m_{1}}\alpha_{n}}\quad\quad\forall n\in\mathbb{N}.$
By the condition 3.4, there exists $M>0$ such that for every $n\in\mathbb{N}$ and $j\leq n$
$\alpha_{n}\leq M(\alpha_{n-j}+\alpha_{j}).$
By using the condition 4.4, we can write that for every $k\in\mathbb{N}$, there exist $m_{2}\in\mathbb{N}$ and $C_{2}>0$ such that
$b_{n,k}\leq C_{2}e^{-\frac{1}{m_{2}}\alpha_{n}}$
for every $n\geq k$. Let say $m_{3}=\min(m_{1},m_{2})$. Then we have
$\begin{split}\|\widecheck{T}_{\theta}e_{n}\|_{k}&=\sup_{1\leq j\leq n}|\theta_{n-j}|b_{j,k}\\ &\leq C_{1}C_{2}\sup_{1\leq j\leq n}e^{-\frac{1}{m_{1}}\alpha_{n-j+1}}e^{-\frac{1}{m_{2}}\alpha_{j}}\\ &\leq C_{1}C_{2}\sup_{1\leq j\leq n}e^{-\frac{1}{m_{1}}\alpha_{n-j}}e^{-\frac{1}{m_{2}}\alpha_{j}}\\ &\leq C_{1}C_{2}\sup_{1\leq j\leq n}e^{-\frac{1}{m_{3}}(\alpha_{n-j+1}+\alpha_{j})}\\ &\leq=C_{1}C_{2}e^{-\frac{1}{m_{3}M}\alpha_{n}}=C_{2}\|e_{n}\|_{m_{3}M}\\ \end{split}$
@@
that is, $\widecheck{T}_{\theta}:\Lambda_{1}(\alpha)\to K(b_{n,k})$ is continuous. ∎
@@
Let $\alpha=(\alpha_{n})_{n\in\mathbb{N}}$ be a non-negative increasing sequence with $\lim_{n\to\infty}\alpha_{n}=\infty$ and $K(b_{n,k})$ be a nuclear Köthe space. Assume that $\alpha$ satisfies the condition 3.4. $\widecheck{T}_{\theta}:\Lambda_{1}(\alpha)\to K(b_{n,k})$ is compact if and only if $\theta\in(\Lambda_{1}(\alpha))^{\prime}$ and the following condition holds:
$\quad\quad\exists m\in\mathbb{N}\quad\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad\quad b_{n,k}\leq Ce^{-\frac{1}{m}\alpha_{n}}\quad\quad\forall n\geq k.$ (4.5)
@@
If $\widecheck{T}_{\theta}$ is compact, then $\theta\in(\Lambda_{1}(\alpha))^{\prime}$ from Proposition 4.1 and Lemma 2.1 gives us that there exists a $m\in\mathbb{N}$ such that for all $k\in\mathbb{N}$
$\sup_{n\in\mathbb{N}}{\|\widecheck{T}e_{n}\|_{k}\over\|e_{n}\|_{m}}<+\infty.$
Then there exists a $C_{1}>0$ such that for every $k\in\mathbb{N}$
$\|\widecheck{T}e_{n}\|=\sup_{1\leq j\leq n}|\theta_{j-n}|b_{j,k}\leq C_{1}\|e_{n}\|_{m}=C_{1}e^{-{1\over m}\alpha_{n}}.$
By assuming $\theta_{0}\neq 0$, for every $k\in\mathbb{N}$ we can find a $C>0$ such that
$b_{n,k}\leq Ce^{-{1\over m}\alpha_{n}}$
for all $n\in\mathbb{N}$. Then 4.5 is satisfied. For the converse, assume that $\theta\in(\Lambda_{1}(\alpha))^{\prime}$ and 4.5 holds. By the same calculation of the proof of the Theorem 4.5, we can write that there exists $m\in\mathbb{N}$ such that for all $k\in\mathbb{N}$
@@
Let $\alpha=(\alpha_{n})_{n\in\mathbb{N}}$ be a non-negative increasing sequence with $\lim_{n\to\infty}\alpha_{n}=\infty$. Assume that $\alpha$ satisfies the condition in 3.4 and there exists a non-decreasing function $S:\mathbb{N}\to\mathbb{N}$ such that the following condition hold:
$\quad\quad\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad\quad b_{n,k}\leq Ce^{-\frac{1}{S(k)}\alpha_{n}}\quad\quad\quad\quad\forall n\in\mathbb{N}.$ (4.6)
@@
is S-tame.
@@
Since $\theta\in(\Lambda_{1}(\alpha))^{\prime}$, there exists some $m_{1}\in\mathbb{N}$ and $C_{1}>0$ such that
$\quad\quad\quad\quad\quad\quad\quad\quad|\theta_{n-1}|\leq C_{1}e^{-\frac{1}{m_{1}}\alpha_{n}}\quad\quad\quad\quad\quad\quad\forall n\in\mathbb{N}.$
By using 4.6, we can write that for every $k\in\mathbb{N}$ satisfying $S(k)\geq m_{1}$, there exists a $C_{2},C_{3}>0$ such that
$\begin{split}\|\widecheck{T}_{\theta}e_{n}\|_{k}&=\sup_{1\leq j\leq n}|\theta_{n-j}|b_{j,k}\\ &\leq C_{1}\sup_{1\leq j\leq n}e^{-\frac{1}{m_{1}}\alpha_{n-j+1}}b_{j,k}\\ &\leq C_{2}\sup_{1\leq j\leq n}e^{-\frac{1}{m_{1}}\alpha_{n-j+1}}e^{-\frac{1}{S(k)}\alpha_{j}}\\ &\leq C_{2}e^{-\frac{1}{S(k)}(\alpha_{n-j+1}+\alpha_{n})}\leq C_{3}e^{-\frac{1}{MS(k)}\alpha_{n}}\leq C_{3}\|e_{n}\|_{MS(k)}\end{split}$
@@
$\|\widecheck{T}_{\theta}e_{n}\|_{k}\leq C_{3}\|e_{n}\|_{S(k)},$
that is the family $\widecheck{\mathcal{A}}_{1}$ is $\tilde{S}$-tame, where $\tilde{S}(n)=MS(n)$ for all $n\in\mathbb{N}$ for M in the condition 3.4. ∎

### v3

@@ 4. Operators Defined by an Upper Triangular Toeplitz Matrix
Let $\Lambda_{\infty}(\alpha)$ be a nuclear power series space of infinite type and $K(b_{n,k})$ be a Montel Köthe space. $\widecheck{T}_{\theta}:\Lambda_{\infty}(\alpha)\to K(b_{n,k})$ is compact if and only if $\theta\in(\Lambda_{\infty}(\alpha))^{\prime}$ and the following condition holds:
$\hskip 14.45377pt\exists m\in\mathbb{N}\quad\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad b_{n-s_{0},k}\leq Ce^{m\alpha_{n}}\hskip 28.90755pt\forall n>s_{0}.$ (4.2)
@@
This gives us that there exists a $C>0$ such that
$|\theta_{s_{0}}|b_{n-s_{0},k}\leq\|\widecheck{T}_{\theta}e_{n}\|=\sum^{n}_{j=1}|\theta_{n-j}|b_{j,k}\leq C\|e_{n}\|_{m}=Ce^{m\alpha_{n}}.$
This says that the condition (4.2) is satisfied.
For the converse, assume $\theta\in(\Lambda_{\infty}(\alpha))^{\prime}$ and the condition (4.2) holds. By the same calculation of the proof of the Theorem 4.3, we can write that there exists a $m\in\mathbb{N}$ such that for all $k\in\mathbb{N}$
@@
Let $\Lambda_{\infty}(\alpha)$ be a nuclear power series space of infinite type and $K(b_{n,k})$ be a Köthe space. Assume that there exists a non-decreasing function $S:\mathbb{N}\to\mathbb{N}$ such that the following condition holds
$\hskip 28.90755pt\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad\quad b_{n,k}\leq Ce^{S(k)\alpha_{n}}\hskip 57.81621pt\forall n\in\mathbb{N}.$ (4.3)
@@
The proof is similiar to the proof of Theorem 4.3. Let us assume $\theta\in(\Lambda_{\infty}(\alpha))^{\prime}$. Then there exists some $m_{1}\in\mathbb{N}$ and $C_{1}>0$ such that
[…]
Since $\Lambda_{\infty}(\alpha)$ is nuclear then $\sum^{\infty}_{n=1}e^{-m_{2}\alpha_{n}}$ is convergent for some $m_{2}\in\mathbb{N}$ and using similiar to steps in the proof of Theorem 4.3 we can show that there exists a $D>0$ such that
$\sum^{n}_{j=1}e^{k\alpha_{j}}\leq De^{(k+m_{2})\alpha_{n}}$
for every $k\in\mathbb{N}$. For every $k\in\mathbb{N}$ satisfying $S(k)\geq m_{1}+m_{2}$, there exists $C_{2},C_{3}>0$ such that
$\begin{split}\|\widecheck{T}_{\theta}e_{n}\|_{k}&=\sum^{n}_{j=1}|\theta_{n-j}|b_{j,k}\leq C_{1}\sum^{n}_{j=1}e^{m_{1}\alpha_{n-j+1}}b_{j,k}\\ &\leq C_{2}e^{m_{1}\alpha_{n}}\sum^{n}_{j=1}e^{S(k)\alpha_{j}}\leq C_{3}e^{m_{1}\alpha_{n}}e^{(S(k)+m_{2})\alpha_{n}}\leq C_{3}e^{2S(k)\alpha_{n}}=\|e_{n}\|_{2S(k)}.\end{split}$
@@
$\|\widecheck{T}_{\theta}e_{n}\|_{k}\leq C_{2}\|e_{n}\|_{2S(k)}$
for all $k$ satisfying $S(k)\geq m_{1}+m_{2}$, that is, the family $\widecheck{\mathcal{A}}_{\infty}$ is 2S-tame. ∎
For the converse of Proposition 4.4, we need a stable sequence $\alpha$ when $K(a_{n,k})$ is power series space of finite type $\Lambda_{1}(\alpha)$.
@@
Let $\alpha=\left(\alpha_{n}\right)_{n\in\mathbb{N}}$ be a stable sequence. $\widecheck{T}_{\theta}:\Lambda_{1}(\alpha)\to K(b_{n,k})$ is well-defined and continuous if and only if $\theta\in(\Lambda_{1}(\alpha))^{\prime}$ and the following condition holds
$\hskip 14.45377pt\forall k\in\mathbb{N}\quad\exists m\in\mathbb{N},C>0\quad\quad\quad\quad b_{n,k}\leq Ce^{-\frac{1}{m}\alpha_{n}}\hskip 36.135pt\forall n\in\mathbb{N}.$ (4.4)
@@
Let assume $\theta$ be a sequence and $s_{0}=\min\{t:\theta_{t}\neq 0\}.$ If the operator $\widecheck{T}_{\theta}$ is continuous, Proposition 4.1 gives us that $\theta\in(\Lambda_{1}(\alpha))^{{}^{\prime}}$ and for every $k\in\mathbb{N}$, there exist $m\in\mathbb{N}$, $C>0$ such that
$b_{n,k}\leq Ce^{-\frac{1}{m}\alpha_{n+s_{0}}}\leq Ce^{-\frac{1}{m}\alpha_{n}}$
for every $n\in\mathbb{N}$. This says that (4.4) holds since $\beta$ is an increasing sequence. For the converse, we assume that assume that $\alpha$ is stable, $\theta\in(\Lambda_{1}(\alpha))^{\prime}$ and (4.4) holds. Then there exist some $m_{1}\in\mathbb{N}$ and $C_{1}>0$ such that
[…]
By stability of $\alpha$, we will show that there exists $M>0$ such that for every $n\in\mathbb{N}$ and $j\leq n$ it holds that
$\alpha_{n}\leq M(\alpha_{n-j+1}+\alpha_{j}).$ (4.5)
Since $\alpha$ is stable, there exists a $M>0$ such that $\alpha_{2t}\leq M\alpha_{t}$ for every $t\in\mathbb{N}$. Assume that $n=2t$ or $n=2t+1$ and $1\leq j\leq n$
$\alpha_{n}\leq\alpha_{2t+2}\leq M\alpha_{t+1}\leq M(\alpha_{n-j+1}+\alpha_{j})$
since in this case $t+1\leq j$ or $t+1\leq n-j+1$ and then we have $\alpha_{n-j+1}+\alpha_{j}\geq\alpha_{t+1}$. Therefore the inequality (4.5) is satisfied. By using the condition (4.4), we can write that for every $k\in\mathbb{N}$, there exist $m_{2}\in\mathbb{N}$ and $C_{2}>0$ such that
$b_{n,k}\leq C_{2}e^{-\frac{1}{m_{2}}\alpha_{n+s_{0}}}\leq C_{2}e^{-\frac{1}{m_{2}}\alpha_{n}}$
for every $n\in\mathbb{N}$. Let us say $m_{3}=\min(m_{1},m_{2})$. Then we have
$\begin{split}\hskip 72.26999pt\|\widecheck{T}_{\theta}e_{n}\|_{k}&=\sum^{n}_{j=1}|\theta_{n-j}|b_{j,k}\leq C_{1}C_{2}\sum^{n}_{j=1}e^{-\frac{1}{m_{1}}\alpha_{n-j+1}}e^{-\frac{1}{m_{2}}\alpha_{j}}\\ &\leq C_{1}C_{2}\sum^{n}_{j=1}e^{-\frac{1}{m_{3}}(\alpha_{n-j+1}+\alpha_{j})}\\ &\leq C_{1}C_{2}\sum^{n}_{j=1}e^{-\frac{1}{Mm_{3}}\alpha_{n}}\leq C_{1}C_{2}ne^{-\frac{1}{m_{3}M}\alpha_{n}}\end{split}$
Now we choose a $m_{4}\in\mathbb{N}$ so that $m_{4}>m_{3}M$, then
$\hskip 36.135pt-\frac{1}{m_{3}M}+\frac{1}{m_{4}}<0\hskip 36.135pt\text{and}\hskip 36.135pt\lim_{n\to\infty}ne^{(-\frac{1}{m_{3}M}+\frac{1}{m_{4}})\alpha_{n}}=0.$
Then there exists a $D>0$ such that for all $n\in\mathbb{N}$
$ne^{-\frac{1}{m_{3}M}\alpha_{n}}\leq De^{-\frac{1}{m_{4}}\alpha_{n}}$
and for some $C_{3}>0$ and for all $n\in\mathbb{N}$
$\|\widecheck{T}_{\theta}e_{n}\|_{k}\leq C_{1}C_{2}ne^{-\frac{1}{m_{3}M}\alpha_{n}}\leq C_{3}e^{-\frac{1}{m_{4}}\alpha_{n}}\leq C_{3}\|e_{n}\|_{m_{4}}.$
@@
that is, $\widecheck{T}_{\theta}:\Lambda_{1}(\alpha)\to K(b_{n,k})$ is well-defined and continuous by Proposition 2.2. ∎
@@
Let $\alpha=(\alpha_{n})_{n\in\mathbb{N}}$ be a stable sequence and $K(b_{n,k})$ be a Montel Köthe space. $\widecheck{T}_{\theta}:\Lambda_{1}(\alpha)\to K(b_{n,k})$ is compact if and only if $\theta\in(\Lambda_{1}(\alpha))^{\prime}$ and the following condition holds:
$\hskip 14.45377pt\exists m\in\mathbb{N}\quad\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad b_{n,k}\leq Ce^{-\frac{1}{m}\alpha_{n}}\hskip 36.135pt\forall n\geq k.$ (4.6)
@@
Let assume $\theta$ be a sequence and $s_{0}=\min\{t:\theta_{t}\neq 0\}.$ If the operator $\widehat{T}_{\theta}$ is continuous, then $\theta\in(\Lambda_{1}(\alpha))^{\prime}$ from Proposition 4.1 and Proposition 4.2 says that for every $k\in\mathbb{N}$, there exist $m\in\mathbb{N}$, $C>0$ such that
$b_{n,k}\leq Ce^{-\frac{1}{m}\alpha_{n+s_{0}}}\leq Ce^{-\frac{1}{m}\alpha_{n}}$
for every $n\in\mathbb{N}$. This says that (4.6) holds since $\alpha$ is an increasing sequence.
For the converse, assume that $\theta\in(\Lambda_{1}(\alpha))^{\prime}$ and the condition (4.6) holds. By the same calculation of the proof of the Theorem 4.6, we can write that there exists $m\in\mathbb{N}$ such that for all $k\in\mathbb{N}$
@@
Let $\alpha=(\alpha_{n})_{n\in\mathbb{N}}$ be a stable sequence and $M=\sup_{n\in\mathbb{N}}\frac{\alpha_{2n}}{\alpha_{n}}$. Assume that there exists a non-decreasing function $S:\mathbb{N}\to\mathbb{N}$ such that the following condition hold:
$\hskip 21.68121pt\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad\quad\quad b_{n,k}\leq Ce^{-\frac{1}{S(k)}\alpha_{n}}\hskip 43.36243pt\forall n\in\mathbb{N}.$ (4.7)
@@
is $\tilde{S}$-tame where $\tilde{S}(k)=(M+1)S(k)$ for every $k\in\mathbb{N}$.
@@
The proof is similar to the proof of the Theorem 4.6. Since $\theta\in(\Lambda_{1}(\alpha))^{\prime}$, there exists some $m_{1}\in\mathbb{N}$ and $C_{1}>0$ such that
[…]
By using (4.7), we can write that for every $k\in\mathbb{N}$ satisfying $S(k)\geq m_{1}$, there exists a $C_{2},C_{3},C_{4}>0$ such that
$\begin{split}\|\widecheck{T}_{\theta}e_{n}\|_{k}&=\sum^{n}_{j=1}|\theta_{n-j}|b_{j,k}\leq C_{1}\sum^{n}_{j=1}e^{-\frac{1}{m_{1}}\alpha_{n-j+1}}b_{j,k}\leq C_{2}\sum^{n}_{j=1}e^{-\frac{1}{m_{1}}\alpha_{n-j+1}}e^{-\frac{1}{S(k)}\alpha_{j}}\\ &\leq C_{2}\sum^{n}_{j=1}e^{-\frac{1}{S(k)}(\alpha_{n-j+1}+\alpha_{n})}\leq C_{3}\sum^{n}_{j=1}e^{-\frac{1}{MS(k)}\alpha_{n}}=C_{3}ne^{-\frac{1}{MS(k)}\alpha_{n}}\\ &\leq C_{4}e^{-\frac{1}{(M+1)S(k)}\alpha_{n}}\leq C_{4}\|e_{n}\|_{(M+1)S(k)}\end{split}$
@@
$\|\widecheck{T}_{\theta}e_{n}\|_{k}\leq C_{3}\|e_{n}\|_{(M+1)S(k)}$
that is the family $\widecheck{\mathcal{A}}_{1}$ is $\tilde{S}$-tame, where $\tilde{S}(k)=(M+1)S(k)$ for all $k\in\mathbb{N}$. ∎